OpenStudy (anonymous):

Use Euler's method with step increment h = 0.1 to approximate y(0.3), where y(x) is the solution to the initial value problem y' = y+xy y(0)=1

OpenStudy (anonymous):

y(.1)=y(0)+ y' (0)h y(.1)=y(0)+ y' (0).1 = 1 + (1+0*1).1=1.1 y(.2)= y(.1)+ y'(.1) (.1) 1.1 + (1.1+(.1*1.1)).1=

OpenStudy (anonymous):

so is the answer about 1.37?